Fermi National Laboratory
 

Accelerator Update for February 13, 2006 - February 15, 2006

 

Monday February 13

The day shift began with Operations monitoring store 4639 and the Recycler stash, with the Antiproton Source (Pbar) stacking, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the store, and with NuMI and MiniBooNE taking beam.

Operation turned off the NuMI beam at 10:40 AM in preparation for an access.

Operations began preparing the Recycler to stash antiprotons at 11:09 AM.

Operations began sending beam to the Meson MCenter experiment at 12:10 PM. The experiment’s analyzing magnet (MC7AN1) had earlier developed a LCW leak. The experiment will run without MC7AN1 until shutdown.

A Main Injector (MI) expert began a study period at 12:28 PM.

At 12:35 PM, the Tevatron Electron Lens (TEL) tripped off twice in six minutes. It reset okay.

The MI expert ended her study at 3:26 PM.

The Recycler began preparing to stash at 3:26 PM. The stash ended. Pbar resumed stacking at 4:16 PM. MCenter also resumed taking beam.

Operations reestablished beam to NuMI at 4:54 PM.

MTest resumed taking beam at 6:03 PM.

The Recycler began preparing to stash at 7:21 PM. The stash ended. Pbar resumed stacking around 8:09 PM. MCenter and MTest also resumed taking beam.

At 8:12 PM, Operations reported that an Accumulator RF station (ARF1) had suffered a power amplifier fault when stacking resumed. It reset okay.

The Recycler began preparing to stash at 11:48 PM.

Tuesday February 14

The midnight shift began with Operations monitoring store 4639, with the Recycler stashing, and with NuMI, MiniBooNE, MTest, and MCenter taking beam.

The stash ended. Pbar resumed stacking at 12:34 AM. MCenter and MTest also resumed taking beam.

The Recycler began preparing to stash at 3:27 AM. The stash ended at 3:55 AM. Pbar resumed stacking around 4:07 AM. MCenter and MTest resumed taking beam.

Tevatron (TeV) experts began an end-of-store study at 7:38 AM. The study ended ten minutes later.

Operations terminated store 4639 at 8:05 AM.

At 8:08 AM, the TeV sector D3 wet engine, which had been showing signs of failing all shift, failed. Cryogenic system technicians investigated and reported that the engine needed a cold valve change. They completed the job thirty minutes later.

Operations prepared D0 and CDF for an access into their collision halls at 8:41 AM.

Cryo system techs reported at 8:50 AM that they were going to change out a TeV sector C2 wet engine flywheel.

The Recycler began preparing to stash at 9:40 AM. The stash ended at 10:03 AM. Pbar resumed stacking around.

CDF came out of access at 10:04 AM.

D0 came out of access at 11:13 AM.

Operations put the TeV into shot setup at 12:05 PM.

Operations established store 4642 at 2:39 PM with an initial luminosity of 114.68E30. Pbar resumed stacking. MCenter and MTest resumed taking beam.

The Recycler began preparing to stash at 4:10 PM. The stash ended at 4:24 PM. Pbar resumed stacking around 5:01 PM. MCenter and MTest resumed taking beam.

At 5:23 PM, The TeV suffered sectors E1 and F4 quenches.

Operations began the TeV turn on at 7:38 PM.

The Recycler began preparing to stash at 9:08 PM.

Operations put the TeV through a dry squeeze at 9:17 PM.

The stash ended at 9:41 PM. MCenter and MTest resumed taking beam. Pbar resumed stacking around 9:50 PM.

At 11 PM, NuMI tripped off on an argon gas flow alarm. Operations contacted experts who all agreed that NuMI could continue running. NuMI turned back on around 11:30 PM.

Wednesday February 15

The midnight shift began with Operations monitoring the stash, with Pbar stacking, and with NuMI, MiniBooNE, MTest, and MCenter taking beam.

Operations put the TeV through a dry squeeze at 12:50 AM.

The Recycler began preparing to stash at 2:17 AM. The stash ended at 2:49 AM. Pbar resumed stacking around 3 AM. MCenter and MTest resumed taking beam.

Operations turned off the NuMI beam at 7:05 AM for a scheduled access.

The Recycler began preparing to stash at 7:14 AM. The stash ended. MCenter and MTest resumed taking beam at 7:44 AM.

Operations set up Pbar for a reverse proton study at 7:49 AM.

The Plans for Wednesday

The plans for later this morning are to establish a special low luminosity store made up of one bunch of protons and one bunch of antiprotons (1x1).

 
Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/15/2006    email Fermilab